Attendees: Bruce Glagola, Nena Moonier, Bruno Fieramosca, Wendy VanWingeren, Shane Flood, Dean Wyncott, and Patti Pedergnana
- The Performance Appraisal system is open for entering self-evaluation information.Bruce will find out the date the electronic entries are due.
- The SR and FE Zones A, C, D, E, and F were walked through yesterday.
- The 14-ID-A-P-01 bellows were removed by beamline technicians yesterday with Greg Banks observing.The bellows manufacturer should be shipping the new parts on Monday.
- Sector 15 has a new portable clean room set up downstream of the A station.
- Eric Rod added guarding to the 434 machine shop lathe.We are awaiting an okay from Frank Curran to release the lock out on this machine.
- 7-BM-B will take white beam 8 a.m. on start up day if Health Physics is available.Bruno is scheduling a BCRRT for next week.
- There are several hard-mounted ODH monitors that need repair or replacing.Bruno has placed handheld monitors in these locations.